# A Perspective on AI-Guided Molecular Simulations in VR: Exploring Strategies for Imitation Learning in Hyperdimensional Molecular Systems

### v1

@@ 4.1.1 Ligand/drug binding to protein
One of the most promising applications of IL for iMD-VR is in computer-aided drug design (CADD)[47]. CADD methods are being used within drug development to reduce the financial and temporal costs associated with the discovery, development and analysis of drug candidates. Sabe et al. [36] reported in 2021 that some form of CADD technique had been used in the development pipeline of more than 70 commercialised drugs. Where the target protein structure is known, MD simulations can be used to shortlist candidate molecules for potential bioactivity by calculating protein-ligand complex stability[43]. However, simulating rare events such as ligand binding using MD still remains a challenge.
@@
iMD-VR has been demonstrated as a human-in-the-loop strategy that leverages the human ability to perform spatial tasks to address the problem of the simulation of ligand binding. Ligand binding is akin to 4-dimensional Tetris, where one 3D shape must fit into another. The difficulty is that these shapes are dynamic and flexible, and that they interact with one another in complex ways. Although this is a difficult task to boil down into an algorithm, humans are able to do this naturally using their spatial intuition combined with their motor skills to perform these types of tasks with minimal training. This is exemplified in a study by Deeks et al.[7], who demonstrated the use of iMD-VR for docking ligands to the main protease of the SARS-CoV-2 virus (the virus responsible for the COVID-19 pandemic). The authors found that iMD-VR experts were able to form docked structures that were in agreement with the crystal structures found experimentally. Another notable study by Deeks et al.[8] found that non-experts could also generate accurate structures of protein-ligand complexes. The authors reported that novice iMD-VR users, many of whom were also not experts in ligand binding, could reliably reproduce experimentally-derived docking poses of flexible ligands with only a short amount of training (<40 minutes in VR). This suggests that IL models could be trained effectively using data gathered from both expert and non-expert users, increasing the size of training sets that leverage human intuition to further sample these non-trivial rare events. IL could greatly enhance the use of iMD-VR in the context of ligand-protein binding by learning from the physically relevant trajectories produced by both experts and non-experts to effectively sample the space of possible binding pathways, leading to a better quantitative understanding of the relative energetics of the docking process that naturally influences the effectiveness of a given drug candidate. This could be extended further to protein engineering by training IL models on iMD-VR-generated datasets of users exploring binding pathways for novel proteins.
4.1.2 Material properties investigation
@@
Crossley-Lewis et al. [5] also used iMD-VR to examine the transport of the catalytic promoter methyl $n$-hexanoate through the H-ZSM-5 zeolite. In this system, the researchers used iMD-VR to sample the dynamics of the promoter-zeolite system after the rare event of desorption of methyl $n$-hexanoate from a Brønsted acid site within the zeolite framework, an event that is unlikely to be observed on the timescale of a typical unbiased MD simulation.[5] By applying biasing forces to desorb the promoter molecule and pull it into varied positions in the zeolite framework, the researchers were able to investigate the transport dynamics of methyl $n$-hexanoate on accessible timescales, identifying features of zeolite structure relevant to the dynamics of the molecule.[5] To develop this study further, a more quantitative understanding of both the energetics of desorption and the rate of diffusion of the promoter after desorption would be desirable. Once again, this is where IL could assist iMD-VR in a research context: not only could IL help to determine the optimal pathways for desorption, but it could be used to enhance sampling of the subsequent dynamics after such events. These dynamics are necessary to better approximate quantities of interest such as the diffusion coefficient, a measure of the average promoter diffusion (that greatly influences the catalytic efficiency of the material in question[5]), helping guide the search for effective catalysts and promoters.
@@ 5 Conclusion
By harnessing the spatial reasoning abilities and domain expertise of researchers performing molecular manipulation tasks in immersive VR environments, rich datasets can be generated to train AI agents via imitation learning techniques. This human-in-the-loop approach shows great promise for efficiently exploring vast conformational spaces of molecular systems. Imitation learning methods have been successfully employed in robotics for learning complex manipulation tasks from demonstrations, and can potentially be adapted to learn policies for interacting with and manipulating molecular structures in iMD-VR. The unique 3D interaction data captured from researchers in iMD-VR systems presents an opportunity to encode human intuition and domain knowledge into the decision-making of AI agents. Key challenges to be addressed include defining appropriate reward functions, and enabling agent generalization across diverse molecular systems. Hybrid approaches combining imitation learning with other AI techniques like reinforcement learning may prove advantageous in this pursuit. While there is still significant work to be done, the intersection of imitation learning and interactive molecular dynamics in VR presents a promising frontier for accelerating scientific discovery and innovation.
@@
This project was supported by the Xunta de Galicia (Centro de investigación de Galicia accreditation 2019–2022, ED431G-2019/04) and the European Union (European Regional Development Fund—ERDF), as well as by the European Research Council under the European Union’s Horizon 2020 research and innovation program through consolidator grant NANOVR 866559.

### v2

@@ 5.1 Ligand/drug binding to proteins
One potential application of IL for iMD-VR is in computer-aided drug design (CADD) [2]. CADD methods are being used within drug development to increase the efficiency of discovery, development and analysis of drug candidates [64]. Where the target protein structure is known, MD simulations can be used to shortlist candidate molecules for potential bioactivity by calculating protein-ligand complex stability [65]. However, simulating ligand binding using MD remains a challenge due to the long simulation timescales required to simulate these rare events, and the associated computational cost.
@@
iMD-VR provides a tool to address the challenge of simulating ligand binding: by harnessing the innate human ability to perform spatial tasks, iMD-VR enables researchers to induce complex binding events on accessible timescales. For example, Deeks et al.[5] demonstrated the use of iMD-VR for docking ligands to the main protease of the SARS-CoV-2 virus (the virus responsible for the COVID-19 pandemic). The authors found that iMD-VR experts were able to form docked structures that were in agreement with the crystal structures found experimentally. Another notable study by Deeks et al.[4] found that non-experts could also generate accurate structures of protein-ligand complexes. The results of these studies suggest that IL models could be trained effectively using data gathered from both expert and non-expert users, increasing the size of training sets that leverage human intuition to further sample these non-trivial rare events. Additionally, if trained on physically relevant trajectories, IL could be used to efficiently sample the space of relevant binding pathways, further enhancing the use of iMD-VR in the context of protein-ligand binding.
5.2 Material properties investigation
@@
Crossley-Lewis et al. [27] also used iMD-VR to examine the transport of the catalytic promoter methyl $n$-hexanoate through the H-ZSM-5 zeolite. Here, the researchers used iMD-VR to sample the dynamics of the promoter-zeolite system after the rare event of desorption of methyl $n$-hexanoate from a Brønsted acid site within the zeolite framework. By applying biasing forces to desorb the promoter molecule and guide it into various positions in the zeolite framework, the researchers were able to (a) investigate the transport dynamics of methyl $n$-hexanoate on accessible timescales, and (b) identify features of the zeolite structure relevant to the dynamics of the molecule. This study could be built upon by performing further quantitative analysis of both the energetics of desorption and the rate of diffusion of the promoter after desorption. An interesting research question is whether IL could enhance the sampling of relevant desorption pathways and desorbed promoter conformations. If successful, this protocol could be integrated into a typical MD workflow to achieve more efficient approximation of quantities of interest such as the diffusion coefficient—a measure of the average promoter diffusion, which greatly influences the catalytic efficiency of the material in question—helping to guide the search for effective catalysts and promoters.
5.3 Citizen science for dataset generation
In order for an ML model to perform well in a defined domain, it must be trained using high quality data. iMD-VR has previously been used to collect data via ‘citizen science’ approaches, in which data collection is outsourced to non-expert users. Deeks et al. [4] reported that novice iMD-VR users, many of whom were also not experts in ligand binding, could reliably reproduce experimentally-derived docking poses of flexible ligands with only a short amount of training (<40<40 minutes in VR). In another study, Shannon et al. [7] demonstrated that non-experts were able to explore the chemical reactivity of propyne with an OH radical, successfully identifying almost all of the important reaction pathways available to the system. Both of these studies indicate that in certain contexts non-chemist citizen scientists can generate chemically relevant datasets. This suggests that data collection need not be limited to experts in computational chemistry and/or iMD-VR. Furthermore, gamification techniques could be employed for data collection to increase participant engagement and reduce the need for bespoke in-person training. Such techniques have already been successfully applied in iMD-VR by Shannon et al. [7] (see above) and Roebuck Williams et al. [67], who showed that non-expert citizen scientists could sense the properties of interactive molecular simulations in VR.
@@ 6 Conclusion
This work has explored the potential of training AI agents using imitation learning on the rich dataset generated from iMD-VR. The immersive nature of VR environments enables researchers to apply their spatial reasoning abilities and domain expertise directly to molecular manipulation tasks, creating valuable demonstration data that captures human intuition and problem-solving strategies. Our proof-of-principle implementing a CNN BC model shows that the data structures in NanoVer can be used to train IL models capable of learning molecular interactions and solving molecular tasks. Drawing parallels from robotics, where IL has proven successful in teaching complex manipulation tasks, we believe that similar techniques can be adapted for molecular systems. Specifically, BC and IRL approaches could be particularly relevant for learning policies that guide molecular interactions in iMD-VR. The unique three-dimensional interaction data captured from expert demonstrations in iMD-VR systems offers unprecedented opportunities to encode domain knowledge into AI decision-making processes. However, several technical challenges remain, including the ensurance of generalization across diverse molecular systems, and handling the high dimensionality of conformational spaces. Future research directions should explore hybrid approaches that combine IL with reinforcement learning techniques, potentially leveraging recent advances in multi-agent systems and adversarial training methods. The integration of these approaches could lead to more robust and adaptable AI agents capable of assisting researchers in exploring vast conformational spaces, ultimately accelerating drug discovery, protein engineering, and materials design processes.
Declarations
@@
Competing interests On behalf of all authors, the corresponding author states that there is no conflict of interest.
[…]
Funding Information This work has been supported by the European Research Council under the European Union’s Horizon 2020 research and innovation programme through consolidator Grant NANOVR 866559, the Xunta de Galicia (Centro de investigación de Galicia accreditation), and the European Union (European Regional Development Fund—ERDF). DRG also thanks the Axencia Galega de Innovación for funding as an Investigador Distinguido through the Oportunius Program.
[…]
Author contribution All authors contributed to the study conception and design. Material and software preparation was done by J.B. Data collection, model development, and results analysis were performed by M.D. The first draft of the manuscript was written by M.D., R.R.W, and H.J.S. All authors reviewed, commented, and provided feedback on all versions of the manuscript. D.R.G. conceptualized the project, contributed to the final version of the manuscript, and supervised the project. All authors read and approved the final manuscript.
[…]
Data Availability Statement The data that support the findings of this study will be available from the corresponding author upon reasonable request.
[…]
Research Involving Human and/or Animals Not Applicable.
[…]
Informed Consent Not Applicable